Selected information regarding Beta Ltd is given below:

  • Cash flow from operations = $1,822,970
  • Depreciation = 177,370
  • Interest expense = $148,375
  • Fixed capital investment = $451,620
  • Working capital investment = $237,280
  • Net borrowing = $328,150

Given a tax rate of 40%, free cash flow to the firm is closest to:

Select one:

a. $1,637,745

b. $1,371,375

c. $1,460,375

Question 4

Question text

Selected information regarding Sentino Investments is given below:

  • Earnings before interest and tax = $548,950
  • Interest expense = $115,250
  • Capital expenditures = $1,150,285
  • Proceeds from sale of long-term assets = $790,390
  • Depreciation expense = $280,355
  • Current assets (excluding cash) = $595,650
  • Current liabilities (excluding short-term debt) = $291,250

Given a tax rate of 40%, free cash flow to the firm is closest to:

Select one:

a. $165,010

b. -$54,570

c. -$222,963

Question 5

Question text

Selected information regarding Tempra Capital is given below:

  • Earnings before interest and tax = $950,125
  • Interest expense = $215,250
  • Fixed capital expenditures = $775,280
  • Depreciation expense = $290,155
  • Working capital investment = $321,255
  • Net borrowing = $540,290

Given a tax rate of 40%, free cash flow available to holders of common equity is closest to:

Select one:

a. $554,885

b. $174,835

c. $742
Answer #1

Question 1

FCFF = CFO + Int(1=tax) - FCInv

FCFF = 1,822,970 + 148,375(1-0.4) - 451,620 = 1,460,375

C

Question 2

FCFF = EBIT(1-t) + Dep - FCInv - WCInv

FCFF = 548,950(1-.04) + 280,355 - (1,150,285 + 790,390)- (Current Assets - Current Liabilities)

FCFF = 548,950(1-.04) + 280,355 - 359,895 - (595,650 - 291,250)

= -54,570

B

FCFE = (EBIT-Int)(1-t) + Dep - FCInv - WCInv + net borrowing

FCFE = 734875(0.6) + 290,155 - 775,280 -321,255 + 540,290

= 174835

B